# A GNN-based Spectral Filtering Mechanism for Imbalance Classification in Network Digital Twin

### v1

@@ 4.3 Multi-class classification results
Table 6: Classification Results for the training in a Network Digital Twin Environment (Domain A) and Real Network Data (Domain C)
@@
The classification reports provide the precision, recall, and F1-score for every class in the Domain A and Domain C datasets, offering a comprehensive performance analysis across domains. In Table 6, the performance on the training data from the network digital twin environment in Domain A is slightly lower compared to the real network environments Domain C, with an average F1-score of 0.93. Precision and recall values are also marginally lower, around 0.94, indicating that the model performs well but may struggle to identify certain classes accurately. Precision measures the proportion of TP predictions out of all positive predictions (TP and FP) made by the model. In the Domain A dataset, Classes 1, 3, 6, 8, 9, 12, and 14 achieved perfect precision (1.00), indicating that the model correctly identified all instances belonging to these classes without any false positives. However, Classes 5, 7, and 13 have relatively lower precision values, suggesting that the model’s predictions for these classes include more false positives. In the Domain C dataset, as shown in Figure 4, most classes achieved high precision values, with Classes 1, 2, 3, 4, 6, 7, 8, 9, 10, 12, 13, 14, and 15 achieving perfect precision (1.00). However, Class 0 has a lower precision value, indicating that the model’s predictions for this class include more false positives.
@@
Recall, also known as sensitivity, measures the proportion of true positive (TP) predictions out of all actual positive instances (TP and false negatives, FN) in the dataset. In Figure 5, classes 1, 3, 6, 8, 9, 12, 14, and 15 achieved perfect recall (1.00), indicating that the model correctly identified all instances of these classes. However, classes 0, 2, 4, 5, 7, 11, and 13 have lower recall values, suggesting that the model missed some instances of these classes. The F1-score, which is the harmonic mean of the precision and recall, provides a balanced measure of the two metrics. This is particularly useful for datasets with class imbalance. The weighted average F1-score for the Domain A dataset is 0.93, indicating good overall performance across all classes. The weighted average F1-score for the Domain C dataset is 0.97, reflecting excellent performance across all classes.
Overall, both datasets demostrate strong performance in terms of precision, recall, and F1-score results for most classes, indicating that the models effectively classify instances across different categories. The results suggest that the model accurately identifies and classifies the majority of data points. However, Class 0 has the lowest recall (0.63) compared to the others, meaning the model might be missing a significant portion of actual Class 0 instances, misclassifying them into other categories. Similarly, Class 5 has a lower recall (0.68) and F1 score (0.81), and class 13 has a slightly lower recall (0.95) compared to others. Further investigation into Classes 0, 5, and 13 is recommended to improve the classification of these classes and enhance the overall model effectiveness of the data.
Figure 6: ROC Curve analysis obtained from a network of digital twin
Figure 7: ROC curve analysis obtained from real network (Domain C)
The confusion matrices in Figure 4 and 5 illustrate the training performance of each failure class in Domains A and Domain C. A high precision indicates that the model makes fewer false positive predictions. In Figure 4, precision varies across classes, with some classes achieving high precision (e.g., Classes 0, 8, 9, 12, 14, 15) and others having lower precision (e.g., Classes 5, 7, 11, 13). Similarly, in Figure 5, precision varies across classes, but overall, it remains high for most classes.
However, the precision, recall, and F1-score remain consistently high at 97% for the real network dataset, demonstrating robust performance. In contrast, the performance on the network digital twin dataset (Domain A) is slightly lower than in Domain C, with an accuracy of 94.27%. The precision, recall, and F1-score also show a slight decline, averaging 94%, indicating that the model in the simulated environment (Domain A) struggled to classify certain failure types.
@@ 5 Conclusion and Future work
In this paper, we propose a novel CF-GNN model as a solution to improve imbalanced classification for 5G network digital twins. We introduce an adaptive spectral filtering method specifically designed for multi-class classification datasets. A critical analysis of the literature highlights the challenges in training datasets, which often lead to class overlapping, particularly in relation to the imbalance ratio. This issue significantly complicates the identification of minority classes, especially as the number of classes increases. Experiment with the state-of-the-art GNN classification techniques have been conducted on real-network and network digital twin datasets to test the effectiveness of our proposed method in filtering the minority classes.
Future work shall continue to expand on our analysis of the effects of multiclass data difficulty for node classification. It might be investigated how the failure classes can be categorized into sub-types and whether there are small clusters of uncommon minority classes. We also intend to introduce efficient GNN designs data level-based node classification models for class-imbalanced graph data to improve classification performance.
6 CRediT authorship contribution statement
Abubakar Isah: Conceptualization, Writing-original draft, Formal analysis, Software, Data curation. Ibrahim Aliyu: Methodology and Review-writing. Sulaiman Muhammad Rashid: Writing-review and editing. Jaehyung Park: Supervision, Review and Editing. Minsoo Hahn: Methodology, Writing and Review. Junsul Kim: Writing-review and editing, Supervision, Conceptualization
@@ 7 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have influenced the work reported in this study.
8 Data availability
Data will be made available on request.

### v2

@@ 4.3 Multi-class classification results
The classification reports provide the precision, recall, and F1-score for every class in the datasets in Domain A and Domain C comprehensively analyzing performance across domains. Table 7 presents slightly lower performance on the training data from the network digital twin environment in Domain A is slightly lower compared to the real network environments in Domain C, with an average F1-score of 0.93. Precision and recall values are also marginally lower, around 0.94, indicating that the model performs well but may not identify some classes accurately. Precision measures the proportion of TP predictions out of all positive predictions (TPs and FPs) made by the model. In the Domain A dataset, Classes 1, 3, 6, 8, 9, 12, and 14 achieved perfect precision (1.00), indicating that the model correctly identified all instances belonging to these classes without any false positives. However, Classes 5, 7, and 13 have lower precision values, suggesting that the model predictions for these classes include increased FPs. In the Domain C dataset (Figure 7), most classes achieved high precision values, with Classes 1-4, 6-10, and 12-15 achieving perfect precision (1.00). However, Class 0 has a lower precision value, indicating that the model predictions include increased FPs.
Table 7: Classification results for training in a network digital twin Environment (Domain A) and real-world network data (Domain C)
@@
Recall, also known as sensitivity, measures the proportion of TP predictions out of all actual positive instances (TPs and FNs) in the dataset. In Figure 8, classes 1, 3, 6, 8, 9, 12, 14, and 15 achieved perfect recall (1.00), indicating that the model correctly identified all instances of these classes. However, Classes 0, 2, 4, 5, 7, 11, and 13 have lower recall values, suggesting that the model missed some instances of these classes. The F1-score, the harmonic mean of the precision and recall, provides a balanced measure of the two metrics, which is helpful for datasets with class imbalance. The weighted average F1-score for the Domain A dataset is 0.93, indicating good overall performance across all classes. The weighted-average F1-score for the Domain C dataset is 0.97, reflecting excellent performance across all classes.
Figure 9: Receiver operating characteristic (ROC) curve analysis of the network of digital twins (Domain A)
Figure 10: Receiver operating characteristic (ROC) curve analysis of the real network (Domain C)
Overall, Figure 10 and Figure 9 demonstrate robust performance in terms of precision, recall, and F1-score results for most classes, indicating that the models effectively classify instances across categories. The results suggest that the model accurately identifies and classifies most data points. However, Class 0 has the lowest recall (0.63) compared to the others, meaning the model might be missing a sizable portion of actual Class 0 instances, misclassifying them as other categories. Similarly, Class 5 has a lower recall (0.68) and F1 score (0.81), and Class 13 has a slightly lower recall (0.95) than others. Further investigation into Classes 0, 5, and 13 is recommended to improve the classification of these classes and enhance the overall model effectiveness.
@@ 5 Conclusion and Future work
This paper proposes the CF-GNN model to address class imbalances in node classification, particularly in 5G network digital twin datasets. The proposed technique integrates a spectral filtering method for multiclass classification, ensuring a more balanced learning process. The critical analysis of the literature identified critical challenges related to dataset imbalances, where majority-class dominance causes suboptimal embeddings for minority-class nodes. This imbalance complicates classification, particularly as the number of classes increases. We examined how loss function consistency varies across different subgraphs, demonstrating that incorporating class frequency and the node degree into the loss formulation improves the distribution balance while preserving the graph structure. Furthermore, we analyzed how nodes contribute to learning class-specific filter parameters. The findings suggest that some nodes play a more significant role in updating filter parameters, influencing the ability of the model to differentiate between classes effectively.
Experimental results on a real network and digital twin datasets validated the effectiveness of the CF-GNN in mitigating class imbalance and enhancing minority-class classification. Future work should further investigate the role of subclass structures within minority groups and refine GNN designs that employ data-level techniques to enhance the classification performance on imbalanced graph data.
@@ 6 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
@@ 7 Acknowledgement
8 Data availability
Data are available at https://aiforgood.itu.int/event/2023-japan-challenge-network-failure-classification-using-network-digital-twin.